OpenStudy (anonymous):

Trini's car breaks down on a highway. Trini estimates that she is 20 to 30 miles from the nearest car repair shop. She calls a towing company that charges a fee of $80 plus $3 per mile to tow a car. If Trini uses this towing company, which is the best estimate for the amount of money, m, she will pay for the company to tow her car? 103 ≤ m ≤ 113 140 ≤ m ≤ 150 140 ≤ m ≤ 170 160 ≤ m ≤ 170

OpenStudy (anonymous):

\[80+20\times 3\leq m\leq 80+30\times 3\] is a start

OpenStudy (anonymous):

trini probably owns a kia

OpenStudy (anonymous):

hahaha very funny